AGw and GCS Commissioning Observing on 05/06 September 2008.

Observers: D. Miller, A. Rakich and J. Hill in Tucson, and J. Storm in Potsdam for the second half
Software: T. Leibold, T. Purdin and S. Hooper for the first 2 hours
Telescope Operator: A. Ceranski

Summary

Made significant progress. In fact this can be considered the first night of Active Optics wake up now that we have solved all the problems of the AGw and IRTC being rotated from June.

Most of our "priority 1" task list has been completed and Jesper also had 2 hours of data taking.

IRTC can now be removed from the telescope and Lucifer put on, but we still need one full night (most likely) with the AGw unit to complete our last tasks, Jesper may need some more time tomorrow night as well, TBA.

Nighttime Activities

02:05 Open dome to clear sky.

02:10 Brief problem with rotator. (HW_FAULT) Was cured by restarting MCSPU software.

02:26 Already in closed loop guiding with GCS! /tmp/guider010424 through 011162

02:32 Sending WFS corrections with GCS! /tmp/wfscimage000146 was first correction of night, and it had ~400 nm rms aberration. Then the wavefront loop converges to 150-200 nm rms. This data set continues through wfscimage000173.

Telescope is having some kind of problem. Some guide images are elongated by 2 arcsec. I've seen a few elevation servo errors up to 0.5 arcsec. Aaron says the wind is calm. (< 2 m/sec). Not yet clear if this is a servo problem or a guiding problem or a vibration problem. The position angle of the dominant error is changing as we track, so it seems to be on the telescope.

03:06 Tweaking up the pointing offsets, IE=-95 CA=-60 (similar to the pointing model default).

03:08 pointto BS_9177 gs=41 Active Guiding and active loops close on the first try.

03:12 increase guiding exposure time to 300 msec on the fly.

03:23 Why does the wavefront aberration convergence off-axis seem to be ~2x what it was on-axis/

Off-axis WFS continues through image 190.

03:24 GCS dies unexpectedly. Seems to be something related to the "stop guiding" at a new preset when two GUIs are running.

03:25 Restart GCS and preset back to on-axis star.

First wfs image was image 191 with larger correction (few hundred) than what we expected.

03:37 Last guide image was guider011682.

Checking IRTC focus

03:38 pointto M5_0738 gs=5 Active (a brighter star so we can study it on the IRTC) Guide star acquisition failed.

03:41 pointto M5_0738 gs=11 Active (star too far off-axis)

03:43 pointto M5_0738 gs=1 Active

03:50 pointto

/Repository/IRTC_Data/080906/

Currently (default) stage is at 17.76 mm in the off-axis position.

focus00036_cube

03:52 20 mm = focus00037_cube (suspect, because we didn't wait for WFS update)

03:54 20 mm = focus00038_cube

03:56 Cube of 6000 x 0.02 sec images for vibration analysis, but a problem with too many frames. ??

04:10 Had to kill and restart our to-station version of irs to unhang IRTC. Didn't have to restart IRTC. We can't reboot IRTC from the downtown to-station since somebody's X intercepts the Cntrl-Alt-Del needed to bring up the task manager.

04:14 1000 frames x 0.002 sec = vibrate_00001_cube.fit @ EL=72 deg, AZ=348

04:18 3000 frames x 0.002 sec = vibrate_00002_cube.fit

04:20 1000 frames x 0.002 sec =vibrate_00003_cube.fit

04:32 wfscimage 254 start changing stage focus again

stage focus = 24

4:35 100 frames X 0.01 sec (overexposed) focus_00039_cube.fit

4:36 100 frames X 0.001 sec focus_00040_cube.fit

stage focus = 28

4:42 100 frames X 0.001 sec (ActOpt Correction) focus_00039_cube.fit

4:43 100 frames X 0.001 sec focus_00007_cube.fit

stage focus = 32

4:46 100 frames X 0.001 sec focus_00008_cube.fit

stage focus = 36

4:53 100 frames X 0.002 sec focus_00009_cube.fit

Last images before new presetTelescope

wfscimage 293. guiderimage 13544

5:05 Guess that a delta of 8 mm is best focus on IRTC. Thus, John changed the on-axis focus stage position in the AIP_L.cfg file to 29.0 mm

=> IRTC focus is much sharper!!

5:19 last WFSC image in sequence was 302

5:20 Gone to M5_0835, guiding and WFS on guide star 1. Testing open-loop tracking. Trying IDL script to setfocus zero (to drive an update on the collimation lookup table) and take IRTC images at prescribed intervals, testing now. IRTC images in /repository/IRTCdata/080906

Testing Open-Loop Tracking

5:37 collimating on star 3.6' off-axis, converged to ~120 nm so episode of not covnerging at the start of the night probably bad dome seeing or something. IE CA IRTC star to centre of chip.
IE -100 CA -68
elevation 80 degrees.
collimating before each data set.

1 ms exposure

5:42 taking set of data: track_00010_cube.fits; single FITS file file with header information /Repository/IRTC.20080906.054241.fits, subsequent time stamps matched to subsequent cubes.

5:44 track_00011_cube.fits
5:46 track_00012_cube.fits
5:48 track_00013_cube.fits
5:50 track_00014_cube.fits
5:52 track_00015_cube.fits
5:54 track_00016_cube.fits
5:57 track_00017_cube.fits
5:59 track_00018_cube.fits
6:00 track_00019_cube.fits
6:02 track_00020_cube.fits
6:04 track_00021_cube.fits

6:14 Repeating now at lower elevation M5_0713 elevation 53 deg. guide star 2, 192" off-axis first wfsc 309, last 318, now at 51 elevation, sending a TRACK preset

2ms exposure

6:23 taking open loop tracking data 6:27 track_00022_cube.fits; single FITS file file with header information /Repository/IRTC.20080906.062722.fits, subsequent time stamps matched to subsequent cubes.

6:29 track_00023_cube.fits
6:31 track_00024_cube.fits
6:33 track_00025_cube.fits
6:35 track_00026_cube.fits
6:37 track_00027_cube.fits
6:39 track_00028_cube.fits
6:41 track_00029_cube.fits
6:43 track_00030_cube.fits
6:45 track_00031_cube.fits
6:47 track_00032_cube.fits

6:48 going back to M5_0713 again, getting EF coefficients for Zernikes Z12-Z22 collimating M1 between every iteration...

7:13 taking 300 0.1 s images (giving 30s) for pre Z12 unaberrated.
7:15 in repository/IRTCdata/080906
7:15 intrafocal Z axis secondary -0.8 mm EF_00033_cube.fit extrafocal +0.8 mm Z-axis secondary EF_00034_cube.fit
7:18 Z12 +500 nm extrafocal EF_00035_cube.fit intrafocal Z-axis secondary EF_00036_cube.fit
7:21 removed Z12, add 500 nm Z13, intrafocal EF_00037_cube.fit extrafocal EF_00038_cube.fit
7:24 removed Z13, back to zero aberration extrafocal EF_00039_cube.fit, intrafocal EF_00040_cube.fit
7:28 add Z14 +500 nm intrafocal EF_00041_cube.fit, extrafocal EF_00042_cube.fit
7:31 remove Z14, add z15 +500, extrafocal EF_00043_cube.fit, intrafocal EF_00044_cube.fit
7:34 remove Z15 +500 nm , back to "unaberrated" intrafocal EF_00045_cube.fit, extrafocal EF_00046_cube.fit
7:37 add Z16 +500, extrafocal EF_00047_cube.fit, intrafocal EF_00048_cube.fit
7:40 remove Z16, add z17 +500, intrafocal EF_00049_cube.fit, extrafocal EF_00050_cube.fit
7:44 remove Z17, unaberrated, extrafocal EF_00051_cube.fit, intrafocal EF_00052_cube.fit
7:47 add Z18 +500, intrafocal EF_00053_cube.fit, extrafocal EF_00054_cube.fit
7:50 remove Z18, add z19 +500, extrafocal EF_00055_cube.fit, intrafocal EF_00056_cube.fit
7:54 remove Z19, unaberrated, intrafocal EF_00057_cube.fit, extrafocal EF_00058_cube.fit
7:59 add Z20 +500, extrafocal EF_00059_cube.fit, intrafocal EF_00060_cube.fit
8:02 remove Z20, add Z21 +500, -0.8 intrafocal EF_00061_cube.fit, +0.8 extrafocal EF_00062_cube.fit
8:05 remove Z21, unaberrated, extrafocal EF_00063_cube.fit, intrafocal EF_00064_cube.fit
8:07 add Z22 -500, intrafocal EF_00065_cube.fit, extrafocal EF_00066_cube.fit Note the minus 500 nm of Z22 sent here.

Done with this, handing over to Jesper for AGw critical stuff.

AGw Calibration to Rotator Axis (Jesper and John)

08:17 pointto M5_0046 Active gs=1 to collimate the telescope (0,1 is a binary pair)

PSF set back to Default distribution of collimation.

Missed star, Adjust IE=-100 CA=-80

8:27 Aaron presets to M103 (coords of last night), collimate. setxy -x 0 -y 612.2

PA=0 = guider000001

PA=90 = guider000002

PA=180 = guider000003

PA=270 = guider000004

08:41 ^C and restart oacserver, home motors, reposition probe

PA=270 = guider000006 (000005 was a bit out of focus, but then so is 6) I forgot to resend focus after homing the motors.

PA=180 = guider000007

setfocus -p 29

PA=180 = guider000008

PA=90 = guider000009 (telescope not tracking)

PA=90 = guider000010

PA=0 = guider000011

08:59 pointto M5_0046 Active gs=2 to collimate the telescope (couldn't find star)

09:02 pointto M5_0046 Active gs=3 to collimate the telescope

09:03 restart GCS and both GCSGUIs.

09:04 pointto still doesn't find star, and GCSGUIs don't show acquisition image.

09:05 Aaron preset back to M103.

PA=0 = guider000014

Slew Rotator 0-360 during 80 sec exposure = guider000015

08:17 Restart oacserver, home motors, reposition probe and focus

08:20 Slew rotator 360-0 during 80 sec exposure = guider000016, 17

08:26 Restart oacserver, etc.

08:31 guider000018

08:36 restart oacserver, etc.

08:40 guider000019

08:42 sequence of readGuideCam -e 500 images to focus. M2 Zglobal increased from 0.20 to 0.28.

Jesper has been adjusting the AGw parameters to adjust the probe so that the center of the circle (the rotator axis) in our rotator slew images is on the hotspot.
radial axis: Ooffset=412720 encoder counts
angular axis: homeOffset 41182

08:52 restart oacserver, etc.

08:53 guider000047 = homeOffset 40000

Jesper says this is making no difference, so put homeOffset back to original value of 41142.

Off-axis collimation tests (Andrew)

09:02 pointto BS_9107 gs=0 Active

10:16 Gone to BS9107 Here we collimate off axis then drive on-axis to measure aberration using set xy for the AGw stage.

10:18 going to guide star #35 GCS died unexpectedly (10:19), then we got an RPC error 10:23 got a star and collimating . y stage at 464 mm wfsc 00008-12 pause guiding set xy to y = 612.2 resume guiding measure wavefronts on axis

All that collimation was a waste of time as I set the y stage to 612.2 and had no star. Of course, because there was a large step between acquiring and guiding for the #35 preset. If we want to change from off-axis to on-axis without sending a preset have to try a different approach, so , will update the collimation on M1 before each off-axis measurement, hope that works.

wfsc13-14 collimating on #35 IGNORE THIS FIRST DATA FORGOT TO PUSH COLLIMATE BUTTON

wfsc 15-17 measuring on axis

going to #32

10:44 GCS died unexpectedly after first guide image in guide box

10:45 invalid buffer RPC error

10:46 GCS died unexpectedly

wfsc 18- collimating on #32

going on-axis GCS died twice

determined that running IDL WFS with the GUI running is causing GCS death, . Shut GUI , gone back to #32 start collimating again...

WFSC 23-24 collimating on #32

WFSC 25-27 measuring on-axis going to guide star #86 collimating WFSC 28-31

going on-axis measuring WFSC 32-34

going to guide star #82 collimating WFSC 35-38

going on-axis measuring wfsc 39-41

going to guidestar 18 collimating wfsc 42-45

going back on-axis measuring WFSC 46-48

going to star #19 couldn't find star going to #39 no star there, going to 46 actually not finding stars because approaching keyhole. gone back on-axis tweaked IE, CA try again...

going to star #19 nothing,

11:52 trying #32 again 32 is there. but guiding fails, watch the star walk through the guiding window... resending preset works this time collimating wfsc 49-51

going on-axis measuring not finding guide star on axis now... long gap befor egetting this data, so wfsc 52-54 measuring on-axis a little suspect.

IE and CA reset to -102 an -72going to off-axis star #87 dammit rotator error and out of time, dawn's first flush puts an end to wavefront sensing.

-- JohnHill - 06 Sep 2008
Topic revision: r10 - 17 Sep 2008, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback